# Beyond Linearity and Time-homogeneity: Relational Hyper Event Models with Time-Varying Non-Linear Effects

### v2

@@ 2.2 Effects in Relational Hyperevent Modeling
The contribution function examines how edge-specific covariates influence the event occurrence rate. Although the true multivariate covariate process $X(t,I,J)$ is unobserved, we can observe $p$-dimensional samples through the vector of measured covariates $\bm{x}(t,I,J)$. The vector $\bm{x}(t,I,J)$ may include exogenous covariates, such as attributes of the senders and/or receivers. For example, in a bibliometric hyper-relational graph of authors and papers, one might use aggregate statistics related to authors’ demographics, academic titles, or institutional affiliations as exogenous covariates. In this study, however, we primarily focus on endogenous covariates. In particular, we highlight a key aspect of event history modeling known as subset repetition (Lerner et al., 2025).
Subset repetition allows for flexible evaluation of social structures based on past events, counting interactions between subgroups of participants over time. To define subset repetition, we begin by quantifying the frequency of interactions between two sets of participants using the activity operator:
@@
where $\binom{H}{\rho}$ denotes the set of all subsets of $H$ of size $\rho$, and $\binom{|H|}{\rho}$ is the number of such subsets. In this work, we use the repeated interaction of sender and receiver subgroups — captured through subset repetition — as one of the key components for explaining hyperevent occurrences.
@@
RHEM formulations in the current state of the art typically rely on the linearity assumption, which assumes that the contribution function is time-invariant and linear in the covariate values. A linear model is expressed as $f^{\text{LE}}\left(\bm{x}(t,I,J),t\right)=\bm{\theta}^{\top}\bm{x}(t,I,J),$ where $\bm{\theta}$ is a vector of unknown coefficients expressing the time-invariant linear effect (LE) of covariates $\bm{x}(t,I,J)$. We aim to relax this assumption by allowing each covariate to contribute in a non-linear way, with effects that may also vary over time. In the relational event modeling literature, more flexible specifications — such as time-varying or non-linear effects — have already been explored (Bauer et al., 2022; Boschi et al., 2025; Filippi-Mazzola and Wit, 2024b). Our goal is to extend these approaches to the relational hyperevent modeling framework, allowing covariate effects that are jointly non-linear and time-varying. We first introduce the mathematical formulations for time-varying effects and non-linear effects, followed by the joint time-varying non-linear effect (TVNLE).
@@
For the remainder of this modeling section, we focus on the contribution of a single covariate $x(t,I,J)\in\bm{x}(t,I,J)$ to the log-rate, denoted by $f^{\square}\left(x(t,I,J),t\right)$. For simplicity, we sometimes omit the dependence on $x(t,I,J)$ and $t$, writing $f^{\square}$ instead. The symbol $\square$ indicates whether the contribution is time-varying ($\square=\text{TVE}$), non-linear ($\square=\text{NLE}$), or jointly time-varying and non-linear ($\square=\text{TVNLE}$). Under the additivity assumption, the overall contribution function is expressed as $f\left(\bm{x}(t,I,J),t\right)=\sum_{k=1}^{p}f^{k,\square}\left(x^{k}(t,I,J),t\right)$, where $x^{k}(t,I,J)$ is the $k$-th covariate in the vector $\bm{x}(t,I,J)$, and $f^{k,\square}$ is its corresponding effect function. Each $f^{k,\square}$ is assumed to be a smooth function.
@@
$f^{\text{TVE}}\left(x(t,I,J),t\right)=\alpha(t)\cdot x(t,I,J),\quad\alpha(t)=\sum_{l=1}^{L}\alpha_{l}a_{l}(t),$ (5)
where $\alpha(t)$ is a smooth function of time evaluated at $t$ and representing the TVE of $x(t,I,J)$. In its simplest form, $\alpha$ is represented as a linear combination of $L$ non-linear basis functions of time $a_{l}$, each weighted by a coefficient $\alpha_{l}$. Note that $f^{\text{TVE}}$ is linear with respect to the covariate value. Thus, $f^{\text{TVE}}$ is the product of a smooth function of time evaluated at $t$ and the covariate value at time $t$. We generally refer to smooth function of time $t$ as $f_{t}$.
@@
Here, $f^{\text{NLE}}$ is a smooth function of the covariate evaluated at time $t$. In its simplest form, $f^{\text{NLE}}$ is represented as a linear combination of $Q$ non-linear basis functions $b_{q}$, each weighted by a coefficient $\beta_{q}$. We generally refer to smooth function of covariate $x$ as $f_{x}$.
We aim to introduce a novel type of effect that, to the best of our knowledge, has not yet been explored in either the relational event or relational hyperevent modeling frameworks. Specifically, we propose using tensor products as smooth functions of several variables (Wood, 2017) — particularly time and covariate — to model the joint time-varying non-linear effect of the covariate on event occurrence. The contribution of a covariate with a TVNLE can be expressed as:
@@
by allowing the coefficients $\beta_{q}$ in Equation (6) to vary smoothly in time, expressing them as smooth functions of time as in Equation (5). We generally refer to smooth function of time $t$ and covariate $x$ as $f_{tx}$.
There are several options for the choice of spline basis functions in Equations (5), (6), and (7). In this paper, we focus on thin plate regression splines (TPRS), which are the default choice in the mgcv package in R (Wood, 2017). While a comprehensive comparison of alternative spline types is beyond the scope of this work, we refer interested readers to Wood (2017) for a detailed discussion. Our primary aim is indeed to underline the importance of incorporating flexible effect specifications in the modeling of relational hyperevents.
@@
TPRS are constructed by transforming and truncating the basis functions obtained from a thin plate spline (TPS) smoothing problem (Wood, 2003). Within the TPS bases, two functions span the subspace of completely smooth terms (Wood, 2017): one is constant over the input variable, and the other is linear, perfectly correlated with the input variable itself. When modeling a time-varying effect $\alpha(t)$, as in Equation (5), these basis functions include a constant term of the form $\alpha_{l}a_{l}(t)=\alpha_{l}a_{l}$, where $a_{l}$ is constant in time. This term behaves like a time-invariant coefficient, representing the effect the covariate would have under a standard linear assumption. In the case of a non-linear effect, as in Equation (6), the contribution $f^{\text{NLE}}$ is expressed as a linear combination of basis functions of the covariate, $\beta_{q}b_{q}(x)$. When the basis function $b_{q}(x)$ is linear, the term $\beta_{q}b_{q}(x)=\beta_{q}b_{q}\cdot x$ corresponds to a linear effect, up to a linear transformation of the covariate itself. This structure can also show that both TVE and NLE can be interpreted as restricted versions of TVNLE, defined in Equation (7). If only the constant basis terms in time are active, the effect becomes non-linear in the covariate but time-invariant. Conversely, if only the linear basis terms in the covariate are retained, the effect becomes time-varying but linear in the covariate.
@@ 2.3 Inference Procedures
To implement the estimation procedure based on the data in Equation (1), we use a sampled version of the partial likelihood (Lerner and Lomi, 2020). Specifically, at each event time, we consider only two instances: the observed hyperevent that actually occurred, and a non-hyperevent — an event that did not occur but could have — randomly sampled from those at risk.
@@
where $(I_{m}^{\ast},J_{m}^{\ast})$ denotes the sender and receiver sets of a non-hyperevent sampled at time $t_{m}$. In our empirical application, the non-event is constrained to have the same sender and receiver cardinalities as the observed event, i.e., $(I_{m}^{\ast},J_{m}^{\ast})\in\binom{V^{I}_{t_{m}}}{|I_{m}|}\times\binom{V^{J}_{t_{m}}}{|J_{m}|}$. Here, $\bm{\theta}$ denotes the vector containing all coefficients associated with the basis functions used to represent the smooth covariate effects, or, in the special case of a linear effect, the single coefficient associated with the covariate. Following Boschi et al. (2025), Equation (8) corresponds to the likelihood of a degenerate logistic regression, where the response is fixed (equal to 1), the intercept is omitted, and the predictors are given by the difference in the contribution function evaluated between the observed event and the sampled non-event. As a result, we can fit our model as a Generalized Additive Model (GAM), rather than relying on traditional survival modeling techniques.
To reduce potential overfitting, especially when using smooth functions, we employ a penalized likelihood as the objective function:
$\mathcal{L}^{\bm{\tau}}(\bm{\theta})=\mathcal{L}(\bm{\theta})-\sum_{k=1}^{p}\tau^{k}P^{k}(f^{k,\square}),$ (9)
where each of the $p$ smooth terms is penalized through an associated penalty term $P^{k}$, with the exception of covariates modeled with a linear effect, for which the penalty is set to zero. The penalty term is weighted by the relative smoothing parameter, controlling the trade-off between model fit and smoothness. Smoothing parameters are stored in vector $\bm{\tau}$.
@@
For a univariate smooth effect $f^{\text{NLE}}$, also denoted by $f_{x}$, a common penalty is $P(f_{x})=\displaystyle\int\left(\frac{\partial^{2}f_{x}}{\partial x^{2}}\right)^{2}dx$, which quantifies the wiggliness of the function. As $\tau\to\infty$, the function $f_{x}$ is increasingly constrained toward linearity, ultimately reducing to a straight line. More complex penalty expressions are used in the case of TPRS. However, expressing $f_{x}$ as a linear combination of basis functions, as in Equation (6), allows each penalty $P_{x}(f_{x})$ to be written as a quadratic form in the corresponding coefficient vector.
@@
When dealing with multivariate smooth functions — and in particular, with functions of both time and covariate, as in our main case of interest — we denote the TVNLE effect by $f_{tx}$. The associated penalty can be expressed as:
@@
where $\tau_{t}$ and $\tau_{x}$ control the the tradeoff between wiggliness in different directions and $f_{t|x}$ is $f_{tx}$ considered as function $t$, holding $x$ constant. If we refer to the previously mentioned $P(f_{x})$ used as an example, the penalty term in the multivariate case would be $P(f_{tx})=\displaystyle\int_{t,x}\tau_{t}\left(\frac{\partial^{2}f_{tx}}{\partial t^{2}}\right)^{2}+\tau_{x}\left(\frac{\partial^{2}f_{tx}}{\partial x^{2}}\right)^{2}\,dtdx$ (Wood, 2017). Estimation of $\bm{\tau}$ is usually performed via cross-validation.

### v3

@@ 3.2 Effects in Relational Hyperevent Modeling
The contribution function examines how edge-specific covariates influence the event occurrence rate. Although the true multivariate covariate process $X(t,I,J)$ is unobserved, we can observe $p$-dimensional samples through the vector of measured covariates $\bm{x}(t,I,J)$. The vector $\bm{x}(t,I,J)$ may include exogenous covariates, such as attributes of the senders and/or receivers. For example, in a bibliometric hyper-relational graph of authors and papers, one might use aggregate statistics related to authors’ demographics, academic titles, or institutional affiliations as exogenous covariates. Figure 1 shows that each author has an attribute called Author Score (S), while each paper has a Paper Impact (I). To explain the rate of publication events, we use the average author score of the authors and the total paper impact of the referenced papers. This is an example of an exogenous hyperevent covariate, derived by aggregating information about the members participating in the hyperevent.
The vector $\bm{x}(t,I,J)$ may also include endogenous covariates. In particular, we highlight a key aspect of event history modeling known as subset repetition (lerner2024relational). Subset repetition allows for flexible evaluation of social structures based on past events, counting interactions between subgroups of participants over time. To define subset repetition, we begin by quantifying the frequency of interactions between two sets of participants using the activity operator:
@@
where $\binom{H}{\rho}$ denotes the set of all subsets of $H$ of size $\rho$, and $\binom{|H|}{\rho}$ is the number of such subsets. In this work, we use the repeated interaction of sender and receiver subgroups – captured through subset repetition – as one of the key components for explaining hyperevent occurrences.
@@
RHEM formulations in the current state of the art typically rely on the linearity assumption, which assumes that the contribution function is time-invariant and linear in the covariate values. A linear model is expressed as $f^{\text{LE}}\left(\bm{x}(t,I,J),t\right)=\bm{\theta}^{\top}\bm{x}(t,I,J),$ where $\bm{\theta}$ is a vector of unknown coefficients expressing the time-invariant linear effect (LE) of covariates $\bm{x}(t,I,J)$. Consider the example in Figure 1, where we examine how the average author score and the total referenced papers’ impact influence publication events. Assuming a linear effect for these covariates implies that their influence on the event rate remains constant between 1950 and 2021, i.e., the effect is homogeneous over time. It also imposes monotonicity: if the coefficient is positive, the contribution increases as more impactful papers are added to the reference list and never reaches saturation. We aim to relax this assumption by allowing each covariate to contribute in a non-linear way, with effects that may also vary over time. In the relational event modeling literature, more flexible specifications – such as time-varying or non-linear effects – have already been explored (bauer2022smooth; boschi2023smooth; filippi2024stochastic; kamalabad2023point; uzaheta2024modeling). Our goal is to extend these approaches to the relational hyperevent modeling framework, allowing covariate effects that are jointly non-linear and time-varying. We first introduce the mathematical formulations for time-varying effects and non-linear effects, followed by the jointly time-varying non-linear effect (TVNLE).
@@
For the remainder of this modeling section, we focus on the contribution of a single covariate $x(t,I,J)\in\bm{x}(t,I,J)$ to the log-rate, denoted by $f^{\square}\left(x(t,I,J),t\right)$. For simplicity, we sometimes omit the dependence on $x(t,I,J)$ and $t$, writing $f^{\square}$ instead. The symbol $\square$ indicates whether the contribution is time-varying ($\square=\text{TVE}$), non-linear ($\square=\text{NLE}$), or jointly time-varying and non-linear ($\square=\text{TVNLE}$). Under the additivity assumption, the overall contribution function is expressed as $f\left(\bm{x}(t,I,J),t\right)=\sum_{k=1}^{p}f^{(k),\square}\left(x^{(k)}(t,I,J),t\right)$, where $x^{(k)}(t,I,J)$ is the $k$-th covariate in the vector $\bm{x}(t,I,J)$, and $f^{(k),\square}$ is its corresponding contribution function. Each $f^{(k),\square}$ is assumed to be a smooth function or a tensor smooth.
a)
b)
Figure 2: Two hypothetical effects in the synthetic motivating example. a) The solid curve represents the estimated time‐varying coefficient $\beta(t)$ of the Average Author Score (AAS), entering the model through the linear effect $\beta(t)\,x^{\text{AAS}}$. Together, these determine the contribution to the log-hazard. During the 1950s, authors with a high average score tended to publish less frequently. Toward the late 1980s this pattern reversed, with high-scoring authors publishing more often, although this tendency subsequently weakened. b) The solid curve represents the contribution $f(x^{\text{RPI}})$ of the total Referenced Papers’ Impact (RPI) to the log-hazard. Initially, higher total RPI increases the likelihood of a publication event. However, when the total impact becomes too large, the contribution decreases, indicating a reduced likelihood of publication.
@@
$f^{\text{TVE}}\left(x(t,I,J),t\right)={\gamma(t)}\cdot x(t,I,J),\quad{\gamma(t)}=\sum_{l=1}^{L}\alpha_{l}a_{l}(t),$ (5)
where $\gamma(t)$ is a smooth function of time evaluated at $t$ and representing the TVE of $x(t,I,J)$. In its simplest form, $\gamma(t)$ is represented as a linear combination of $L$ non-linear basis functions of time $a_{l}(t)$}, each weighted by a coefficient $\alpha_{l}$. Note that $f^{\text{TVE}}$ is linear with respect to the covariate value. Thus, $f^{\text{TVE}}$ is the product of a smooth function of time evaluated at $t$ and the covariate value at time $t$. We generally refer to smooth function of time $t$ as $f_{t}$. The scientific community’s perception of author scores may change. Although the score is synthetically generated, it reflects real situations in which features of scientific authors are viewed differently over time, for example due to new tools available to the community. In practice, this Average Author Score (AAS) does not contribute homogeneously to the log-occurrence rate but instead has a time-varying effect. According to Equation (5), this effect is a smooth function $\gamma(t)$}. Its contribution to the log-hazard is obtained by multiplying the value of this function, at each time point, by the corresponding value of the covariate. An estimate of AAS’s effect is shown in Figure 2a. Details of the estimation procedure are provided in Section 3.3.
@@
Here, $f^{\text{NLE}}$ is a smooth function of the covariate evaluated at time $t$. In its simplest form, $f^{\text{NLE}}$ is represented as a linear combination of $Q$ non-linear basis functions $b_{q}$, each weighted by a coefficient $\beta_{q}$. We generally refer to smooth function of covariate $x$ as $f_{x}$. It is plausible to suspect that citing high-impact papers may increase the likelihood of a paper being published. However, this effect may not be linear and could reach saturation. To investigate this, we model the non-linear influence of the total Referenced Papers’ Impact (RPI) on the log-hazard, as shown in Figure 2 b). As it is discussed in Section 3.3, estimated non-linear functions should be interpreted based on their relative changes and trends and not on their sign. The plot suggests that while an increasing total RPI initially raises the probability of publication, this effect reaches a peak, after which citing additional highly impactful papers appears to have diminishing or even counterproductive effects.
a)
b)
Figure 3: Jointly time-varying non-linear effects of covariates in the synthetic motivating example. In both heatmaps, the x-axis represents time; fixing a point on the y-axis allows us to examine how the effect evolves over time. Since the y-axis represents covariate values, fixing a point on the x-axis shows the non-linear contribution of the covariate. Lighter (yellow) areas indicate a larger contribution to the log-hazard than darker (blue) areas. Gray points are values not observed in the data. Contour lines trace regions of constant effect; closer contours indicate faster changes. a) The heatmap shows that the contribution of the Average Author Score to the log-hazard varies non-monotonically over time and across covariate values. For higher values of AAS (approximately 1.7–1.85), the effect first decreases, then increases, and finally weakens again. This reflects the behavior observed in Figure 2 a). At the beginning and end of the observation period (approximately 1950–1975 and 2000–2021), the trend is clearly non-monotonic, peaking around a value of 1.6. Between 1975 and 2000, the trend is inverted, favoring a higher average author score. b) The heatmap shows the contribution of the total Referenced Papers’ Impact. Although the effect is non-linear (peaking around a value of 3), as in Figure 2 b), it remains relatively stable over time. This is visible from the horizontal contour lines: fixing a point on the y-axis reveals little temporal variation.
Shape-constraint non-linear effect.
It is possible to add shape constraints to the non-linear and time-varying effects in the RHEM by allowing the component smooth functions $f^{(k),\square}$ to satisfy monotonicity or convexity constraints. The idea is based on the Shape Constrained Additive Models (SCAM) extension of generalized additive models (pya2015shape). They use a modified spline basis – called SCOP-spline – in which the constraints are enforced through a re-parameterization of the spline coefficients. Consider the spline representation:
$f^{\text{SCNLE}}(x(t,I,J))=\sum_{q=1}^{Q}\delta_{q}B_{q}(x(t,I,J)),$
where $\{B_{q}\}$ are B-spline basis functions. A sufficient condition for $f^{\prime}(x)\geq 0$ is that the coefficients form a non-decreasing sequence, $\delta_{1}\leq\delta_{2}\leq\cdots\leq\delta_{Q}.$ SCAMs impose this by reparameterizing $\bm{\delta}=\Sigma\ \bm{\tilde{\beta}}$, where $\bm{\tilde{\beta}}=(\beta_{1},\exp(\beta_{2}),\ldots,\exp(\beta_{Q}))^{\top},$ and $\Sigma$ is a cumulative-sum matrix ($\Sigma_{lq}=1$ for $l\geq q$ and 0 otherwise). With this parameterization, the constraints are automatically satisfied.
Jointly time-varying non-liner effect.
We aim to introduce a novel type of effect that, to the best of our knowledge, has not yet been explored in either the relational event or relational hyperevent modeling frameworks. Specifically, we propose using tensor products as smooth functions of several variables (wood2017generalized) – particularly time and covariate – to model jointly time-varying non-linear effect of covariates on event occurrence. The contribution of a covariate with a TVNLE can be expressed as:
@@
by allowing the coefficients $\beta_{q}$ in Equation (6) to vary smoothly in time, expressing them as smooth functions of time as in Equation (5). We generally refer to a smooth function of time $t$ and covariate $x$ as $f_{tx}$.
Consider again the two previously introduced effects and their associated covariates. We have seen that the Average Author Score exhibits a time-varying effect, but it may also display a non-monotonic pattern, similar to the total Referenced Papers’ Impact, where increases in the covariate can eventually become counterproductive for publication activity. Likewise, the influence of RPI may itself have evolved over time. To accommodate these possibilities, we relax the assumption of a smooth effect in only one dimension and instead allow for a jointly non-linear time-varying effect. Figure 2 shows the corresponding estimates. These estimates illustrate how the effect varies across both covariate values and time. In the plots, time is displayed on the x-axis; thus, to examine the temporal evolution of the effect, we fix a point on the y-axis. For higher AAS (approximately 1.7-1.85), the pattern in Figure 2 a) becomes evident: the effect first decreases, then increases, and finally weakens again. Conversely, since the y-axis represents covariate values, fixing a point on the x-axis allows us to observe the non-linear contribution of the covariate. At the beginning and end of the observation period (approximately 1950-1975 and 2000-2021), the trend is clearly non-monotonic, reaching a maximum for values around 1.6. Between 1975 and 2000, the trend is inverted, first decreasing and then increasing, favoring higher average author scores. For the AAS, including this jointly time-varying non-linear effect is crucial. In contrast, the total RPI appears largely stable over time. The non-linear pattern shown in Figure 2 b) can be observed by fixing any point on the x-axis and examining the variation across covariate values, showing a peak around a value of 3. However, when fixing a point on the y-axis, the effect remains fairly homogeneous over time.
There are several options for the choice of spline basis functions in Equations (5), (6), and (7). In this paper, we focus on thin plate regression splines (TPRS) and tensor product smooths, both implemented in the mgcv package in R (wood2017generalized). While a comprehensive comparison of alternative spline types is beyond the scope of this work, we refer interested readers to wood2017generalized for a detailed discussion. Our primary aim is indeed to underline the importance of incorporating flexible effect specifications in the modeling of relational hyperevents.
@@
TPRS are constructed by transforming and truncating the basis functions obtained from a thin plate spline (TPS) smoothing problem (wood2003thin). Within the TPS bases, two functions span the subspace of completely smooth terms (wood2017generalized): one is constant over the input variable, and the other is linear, perfectly correlated with the input variable itself. When modeling a time-varying effect $\gamma(t)$}, as in Equation (5), these basis functions include a constant term of the form $\alpha_{l}a_{l}(t)=\alpha_{l}a_{l}$, where $a_{l}$ is constant in time. This term behaves like a time-invariant coefficient, representing the effect the covariate would have under a standard linear assumption. In the case of a non-linear effect, as in Equation (6), the contribution $f^{\text{NLE}}$ is expressed as a linear combination of basis functions of the covariate, $\beta_{q}b_{q}(x)$. When the basis function $b_{q}(x)$ is linear, the term $\beta_{q}b_{q}(x)=\beta_{q}b_{q}\cdot x$ corresponds to a linear effect, up to a linear transformation of the covariate itself. This structure can also show that both TVE and NLE can be interpreted as restricted versions of TVNLE, defined in Equation (7). If only the constant basis terms in time are active, the effect becomes non-linear in the covariate but time-invariant. Conversely, if only the linear basis terms in the covariate are retained, the effect becomes time-varying but linear in the covariate.
@@ 3.3 Inference Procedures
To implement the estimation procedure based on the data in Equation (1), we use a sampled version of the partial likelihood (lerner2020reliability). Specifically, at each event time, we consider only two instances: the observed hyperevent that actually occurred, and a non-hyperevent – an event that did not occur but could have – randomly sampled from those at risk. Consider the first hyperevent in Figure 1. The observed publication event involves authors 2 and 3 citing papers A and B, as also reported in the left panel of Figure 4. This event represents just one of the 45 possible events that could have occurred at that time, given all potential collaborations among authors 1-4 and papers A and B. Other panels of Figure 4 show examples of non-hyperevents.
23ABCa) Observed hyperevent12ABb) Candidate non-hyperevent 11234Ac) Candidate non-hyperevent 23Bd) Candidate non-hyperevent 3
Figure 4: Observed hyperevent and candidate non-hyperevents in the synthetic motivating example. a) Observed hyperevent. This corresponds to the first hyperevent shown in Figure 1 and listed in Table 1, where authors 2 and 3 publish article C citing papers A and B. To compute the likelihood in Equation 8, a non-hyperevent must be sampled for comparison. Panels b), c), and d) illustrate possible candidates. In the empirical application, non-hyperevents are sampled to have the same cardinality as the observed hyperevent. Under this criterion, only the non-hyperevent in panel b) is a valid candidate among those shown.
@@
where $(I_{m}^{\ast},J_{m}^{\ast})$ denotes the sender and receiver sets of a non-hyperevent sampled at time $t_{m}$. In our empirical application, the non-event is constrained to have the same sender and receiver cardinalities as the observed event, i.e., $(I_{m}^{\ast},J_{m}^{\ast})\in\binom{V^{I}_{t_{m}}}{|I_{m}|}\times\binom{V^{J}_{t_{m}}}{|J_{m}|}$. Practically, this means that among the three non-hyperevents shown in Figure 4, only the first one would be eligible. Here, $\bm{\theta}$ denotes the vector containing all coefficients associated with the basis functions used to represent the smooth covariate effects, or, in the special case of a linear effect, the single coefficient associated with the covariate. The idea behind the likelihood in Equation 8 is that, at each time point, an observed hyperevent is compared with a sampled non-hyperevent – similar to comparing panels a) and b) of Figure 4. To assess whether a covariate is relevant, we examine how its value differs between hyperevents that actually occur and their corresponding non-hyperevents. This difference then serves as the explanatory variable in the logistic formulation introduced above. In this context for hyperevent in a) and non-hyperevent in b), we evaluate AAS and the total RPI. Following boschi2023smooth, Equation (8) corresponds to the likelihood of a degenerate logistic regression, where the response is fixed (equal to 1), the intercept is omitted, and the predictors are given by the difference in the contribution function evaluated between the observed event and the sampled non-event. As a result, we can fit our model as a Generalized Linear Model (GLM) with a logit link function, rather than relying on traditional survival modeling techniques. When the linearity assumption is relaxed and replaced with an additivity assumption – as done in Section 3.2 – the model becomes a Generalized Additive Model (GAM). This framework allows for more flexible and complex effects, such as those illustrated in Figures 2 and 3.
To reduce potential overfitting, especially when using smooth functions, we employ a penalized log-likelihood as the objective function. The penalized log-likelihood is defined as the log of Equation (8) minus a penalty term $P$. This term is a weighted sum of individual penalties $P^{(k)}$, each associated with one of the $p$ smooth components. Covariates modeled linearly receive no penalty, i.e., their penalty is zero. Each penalty is weighted by a smoothing parameter, which controls the trade-off between model fit and smoothness. These smoothing parameters are collected in the vector $\bm{\tau}$. Using R package mgcv, smoothing parameters can be estimated directly from the data using cross-validation, restricted maximum likelihood, or maximum likelihood.
@@
For a univariate smooth effect $f^{\text{NLE}}$, also denoted by $f_{x}$, a common penalty is $P(f_{x})=\displaystyle\int\left(\frac{\partial^{2}f_{x}}{\partial x^{2}}\right)^{2}dx$, which quantifies the wiggliness of the function. As $\tau\to\infty$, the function $f_{x}$ is increasingly constrained toward linearity, ultimately reducing to a straight line. More complex penalty expressions are used in the case of TPRS. However, expressing $f_{x}$ as a linear combination of basis functions, as in Equation (6), allows each penalty $P_{x}(f_{x})$ to be written as a quadratic form in the corresponding coefficient vector. To control smoothness, SCAMs use a difference penalty on the (unconstrained) spline coefficients $\bm{\beta}$, $P(\bm{\beta})=\nu\lVert D\bm{\beta}\rVert^{2},$ where $D$ is a first- or second-order finite-difference operator.
@@
When dealing with multivariate smooth functions – and in particular, with functions of both time and covariate, as in our main case of interest – we denote the TVNLE effect by $f_{tx}$. The associated penalty can be expressed as:
@@
where $\tau_{t}$ and $\tau_{x}$ control the the tradeoff between wiggliness in different directions and $f_{t|x}$ is $f_{tx}$ considered as function $t$, holding $x$ constant. If we refer to the previously mentioned $P(f_{x})$ used as an example, the penalty term in the multivariate case would be $P(f_{tx})=\displaystyle\int_{t,x}\tau_{t}\left(\frac{\partial^{2}f_{tx}}{\partial t^{2}}\right)^{2}+\tau_{x}\left(\frac{\partial^{2}f_{tx}}{\partial x^{2}}\right)^{2}\,dtdx$ (wood2017generalized).
In practice, we use the gam function from the mgcv package in R (wood2017generalized), that automatically implements the penalization. To fit a TVE, we include the term s(time, by=delta_x), where time corresponds to the event times and delta_x represents the difference between the covariate evaluated for the event and the non-event. This fits a smooth function of time multiplied by the covariate. An example of the estimate (solid), along with its confidence intervals (dashed), is shown in Figure 2 a). To fit a NLE, we include the term s(X, by=I), where X is a matrix with the first column containing the covariate values for the event and the second column containing those for the non-event. I is a two-column matrix, with the first column filled with $1$s and the second with 1-1s. An example of an estimated NLE (solid), along with its confidence intervals (dashed), is shown in Figure 2 b). The estimated function cannot be interpreted by its sign due to a centering constraint; only its relative change and trend are meaningful. Finally, to estimate a TVNLE, we include the term te(T, X, by=I), where X and I are defined above, and T is a matrix with both columns equal to the vector of event times. As with the NLE, centering constraints are applied. Therefore, the estimated function needs to be centered, so that its average at each time point is zero. Two examples of centered TVNLE are reported in Figure 3, with the gray squares indicating combinations of covariate and time values that not observed in the data. As before, the estimated effect cannot be interpreted by its sign, only by its relative change and trend. On contour lines, the effect remains constant; the closer the contour lines, the faster the effect changes.
Avoiding overfitting.
Even though our relational event model allows for smooth, non-linear effects, the estimation procedure is designed to prevent these flexible components from overfitting the data. As we described abov, each smooth term is equipped with a penalty that determines how freely it can adapt to patterns in the covariates. Rather than setting this by hand, we let the data decide: the inference scheme treats these smooth components as random effects and selects their level of flexibility by maximizing the marginal likelihood. In practical terms, this means that overly complex curves are automatically penalized, while genuinely nonlinear structures are retained. As a result, the estimated model possesses the expressive power of non-linear effect without succumbing to the instability and overfitting often associated with highly flexible specifications.
3.4 Computational Cost
In this section, we have shown how to define and estimate different types of effects – linear, time-varying, non-linear, shape-constraint non-linear and jointly time-varying non-linear. All inference procedures rely on optimizing the penalized log-likelihood of a logistic additive regression. For a model matrix of dimension $D$, the main computational cost arises from inverting the Hessian matrix, which requires $\mathcal{O}(D^{3})$ operations. A covariate with a linear effect uses a single column in the model matrix, while time-varying and non-linear effects use $L$ and $Q$ columns, respectively. A jointly time-varying non-linear effect uses $L\times Q$ elements, with the tensor product smooth for an observation $m$ given by the Kronecker product of the marginal smooth evaluations. When all $p$ covariates are fitted with the same type of effect, the computational cost increases from $\mathcal{O}(p^{3})$ to $\mathcal{O}(p^{3}L^{3}Q^{3})$ operations. Memory requirements follow the same principle, growing from $O(np)$ for linear effects to $O(npLQ)$ for smooth effects. Importantly, $L$ and $Q$ are constants and independent of network features.
The cost for computing hyperedge covariates $x(t,I,J)$ is not affected by the effects becoming non-linear and/or time-varying. lerner2024relational demonstrate that the computation of hyperedge covariates can scale to networks with more than a million nodes and events. Notably, the logistic regression formulation reduces the computational cost of likelihood evaluation, making it linear in the number of events (lerner2020reliability; boschi2023smooth). Importantly, this computation does not depend on the size of the relational network, unlike full partial likelihood-based estimation procedures.